I spent 1/3 of my money on a sweater. I spent 3/5 of the remainder on a pair of jeans. If I have $36 left, how much did the sweater cost?
Answer by macston(5194)

.
x=original amount
.
(1/3)x+(3/5)(x-(1/3)x)+$36=x
(1/3)x+(3/5)(2/3)x+$36=x
(1/3)x+(6/15)x+$36=x
(11/15)x+$36=x
$36=(4/15)x
$135=x You originally had $135.
.
You spent 1/3 on the sweater:
(1/3)($135)=$45
.
ANSWER: The sweater cost $45.